9. Job 593 was recently completed. The following data have been recorded on its job cost sheet:
  

  Direct materials $2,412
  Direct labor-hours 74 labor-hours
  Direct labor wage rate $    21 per labor-hour
  Machine-hours 137 machine-hours

   

The Corporation applies manufacturing overhead on the basis of machine-hours. The predetermined overhead rate is $22 per machine-hour. The total cost that would be recorded on the job cost sheet for Job 593 would be:
Answer #1

Total direct labor=(74 labor hours*$21 per labor hour)

=(74*21)=$1554

Total overhead would be=$22 per machine hour*137 machine hours

=$3014

Hence total cost=Direct materials+Direct labor+Total overhead

=(2412+1554+3014)

=$6980